# LightCode: Light Analytical and Neural Codes for Channels with Feedback

### v2

@@ Abstract
The design of reliable and efficient codes for channels with feedback remains a longstanding challenge in communication theory. While significant improvements have been achieved by leveraging deep learning techniques, neural codes often suffer from high computational costs, a lack of interpretability, and limited practicality in resource-constrained settings.
We focus on designing low-complexity coding schemes that are interpretable and more suitable for communication systems. We advance both analytical and neural codes. First, we demonstrate that PowerBlast, an analytical coding scheme inspired by Schalkwijk-Kailath (SK) and Gallager-Nakibŏglu (GN) schemes, achieves notable reliability improvements over both SK and GN schemes, outperforming neural codes in high signal-to-noise ratio (SNR) regions. Next, to enhance reliability in low-SNR regions, we propose LightCode, a lightweight neural code that achieves state-of-the-art reliability while using a fraction of memory and compute compared to existing deep-learning-based codes. Finally, we systematically analyze the learned codes, establishing connections between LightCode and PowerBlast, identifying components crucial for performance, and providing interpretation aided by linear regression analysis.
@@ I Introduction
Shannon introduced the concept of feedback channel in [1]. In a channel with feedback, the transmitter cooperates with the receiver to improve the probability of successful transmission by utilizing the feedback from the receiver. In [1], Shannon assumes a perfect noiseless feedback channel with unit delay and demonstrates that the availability of feedback at the transmitter does not change the capacity of the resultant forward channel for memoryless channels. Interestingly, while the capacity remains the same, significant improvements in terms of error exponents can be achieved with the help of feedback.
One of the seminal schemes in the noiseless feedback setting has been provided by Shalkwijk and Kailath in [2, 3] (SK) using a simple linear encoding scheme resulting in a doubly exponential error exponent for finite block lengths. In [4], a two-phase variant of the SK scheme, which we refer to as Gallager-Nakibŏglu (GN), was discussed that can lead to further exponential decay of the error, provided a sufficiently good SNR is available for the forward channel. The SK scheme has been enhanced in various ways. The Modulo-SK scheme [5] and schemes by Chance-Love [6] and Mishra et al [7], extends the SK scheme for noisy feedback, while compressed error correction (CEC) [8] and accumulative iterative code (AIC) [9] focus on reducing channel use through continuous error vector compression, using noiseless feedback.
@@
Recent advances in deep learning revived the interest in coding for channels with feedback by leveraging the expressive power [12] of deep neural networks. Several works proposed deep learning approaches to improve the performance of channel codes, ranging from augmenting analytical decoders with learnable parameters [13, 14, 15, 16] to creating novel neural network architectures based neural encoders and decoders [17, 18] and improving the code design using sequential models [19, 20]. For channels with feedback, deep learning techniques were used to design new encoding and decoding schemes that take advantage of the high-capacity feedback channel. Deepcode [21] modeled both encoder and decoder functions as recurrent neural networks (RNNs) to process a bit stream in a sequential manner to directly minimize the end-to-end transmission error over an AWGN channel. Several works further explored this idea of using learning-based approaches for modeling the encoding and decoding operations, which can be broadly classified into the RNN family of codes [22, 23, 24] and the transformer family of codes [25, 26], discussed in detail in Section V. The current state-of-the-art is generalized block attention feedback (GBAF) [26], which uses self-attention and transformer architecture to perform block coding.
While these state-of-the-art deep-learning-based feedback codes provide tremendous improvements in BLER performance, they also come with significant memory and computational costs that may not be supported by the next generation of communication transceivers that operate with limited onboard resources. Therefore, it is desirable to develop lightweight codes that use simple schemes while providing desirable performance. To accomplish a reduced complexity coding scheme, we explore two different directions in this work. First, we review and understand the existing analytical feedback coding schemes to identify the limitations and propose a new feedback coding scheme that provides non-trivial performance improvements over the existing schemes for channels with passive, noiseless feedback. Next, we propose a lightweight deep-learning-based feedback coding scheme that can greatly reduce the complexity compared to neural block-feedback coding schemes by limiting ourselves to a symbol-by-symbol scheme instead of block coding schemes.
@@
  • We propose PowerBlast, an analytical coding scheme that noticeably improves the performance over Schalkwijk-Kailath and two-phase Gallager-Nakibŏglu schemes (Section IV) and exhibits reliability comparable to state-of-the-art deep-learning-based coding schemes in regions of high-SNR (Section VII).
@@
  • We analyze the representations learned by the LightCode encoder using linear regression and draw comparisons with PowerBlast. We also demonstrate that the relation between encoded symbols and the feedback is highly non-linear in the final rounds, underscoring the need for deep-learning-based coding scheme (Section VIII).
@@ II System Model
We consider a feedback coding scheme with an AWGN forward channel and an AWGN feedback channel with noisy passive feedback. The goal is to transmit a message vector $\mathbf{u}\in\{0,1\}^{K}$ of length $K$ to the receiver using a total of $D$ channel uses, resulting in an overall coding rate of $R=\nicefrac{{K}}{{D}}$. In this work, we consider symbol-by-symbol coding, where the entire block of $K$ bits will be mapped to one symbol. Hence, the terms block and symbol can be used interchangeably. Based on this, the number of rounds of communication for a rate $\nicefrac{{K}}{{D}}$ code is $D$.
@@
where $n_{1}\sim\mathcal{N}(0,\sigma^{2}_{ff})\in\mathbb{R}$ denotes the feedforward noise. The receiver then sends the noisy received symbol as feedback using the feedback channel as
@@
where $\hat{\mathbf{u}}\in\{0,1\}^{K}$. The objective is to design an encoder-decoder pair $\{\phi,\psi\}$ that minimizes the probability of error $\mathbb{P}\{\mathbf{u}\neq\hat{\mathbf{u}}\}$ for a given number of rounds $D$, under a sum power constraint of $\sum_{i=1}^{D}\mathbb{E}[|x_{i}|^{2}]=D$.
@@ III Analytical Coding Schemes for Channels with Noiseless Feedback
In this section, we review analytical coding schemes for channels with noiseless feedback i.e., $\sigma_{fb}^{2}=0$. We begin by reviewing the celebrated Schalkwijk-Kailath coding scheme [2], one of the seminal works in coding for channels with noiseless feedback. We then proceed to review a less widely-known scheme by Gallager-Nakibŏglu [4]. This scheme is similar to the SK scheme for the first $D-1$ rounds. Still, it deviates greatly in the last round, tailoring for the transmission of discrete messages, often significantly improving the performance in one round of communication. The Gallager-Nakibŏglu (GN) scheme is typically not considered as a baseline as it exhibits a worse error rate compared to SK. However, in the next section, we introduce a novel analytical coding scheme, building on the SK and GN schemes. This new scheme achieves a significantly lower error rate than either the SK or GN schemes and is often comparable to highly complex neural coding schemes in high-SNR regions, defying the conventional belief that neural coding schemes are much more reliable than analytical ones.
@@ III-A Schalkwijk-Kailath coding scheme
SK scheme considers the problem of transmitting a fixed number of bits on an AWGN forward channel and a noiseless feedback channel. The transmission begins by mapping $K$ bits of information symbols to a single $2^{K}$-ary pulse-amplitude modulation (PAM) symbol $\Theta$ from the constellation
$\Theta\in\{\pm 1\eta,\pm 2\eta,\dots,\pm(2^{K}-1)\eta\},$ (6)
where $\eta=\sqrt{\nicefrac{{3}}{{(2^{2K}-1}})}$ is the scaling factor to ensure unit power normalization of the PAM constellation. We note that even though a block of information is transmitted since all the $K$ bits of information are mapped to one PAM symbol, the SK scheme at its core can be considered a symbol-by-symbol feedback scheme. We now describe the SK coding scheme in detail.
@@
where $n_{1}\in\mathcal{N}(0,\sigma_{ff}^{2})$ is the noise in the forward AWGN channel. The received symbol $y_{1}$ is then sent back to the transmitter noiselessly. For the second round, the transmitter first computes the receiver’s estimate of the transmitted symbol based on $y_{1}$ as $\hat{\Theta}_{1}=\frac{y_{1}}{1+\sigma^{2}/P}$ using linear minimum mean square error (LMMSE) estimator. In the second round, after receiving $y_{1}$ as feedback, the transmitter sends the scaled version of the error in the estimate from the previous round, i.e., $\epsilon_{1}=\hat{\Theta}_{1}-\Theta$. This process continues for the remaining rounds. In other words, starting from round $i=2$, the goal of the transmitter is to transmit the error in estimate from the previous round, $\epsilon_{i-1}=\hat{\Theta}_{i-1}-\Theta$, after scaling appropriately to satisfy the power constraint. The complete algorithm is described in Alg. 1.
@@
where $S$ is the SNR of the forward AWGN channel on a linear scale.
Input: Message symbol $\Theta$, number of rounds $D$, noise variance $\sigma^{2}$
Round 1: Tx: Power normalization: $x_{1}=\sqrt{P}\Theta$;
Forward channel: $y_{1}=x_{1}+n_{1}$;
Rx: LMMSE estimate of transmit symbol $\hat{\Theta}_{1}=\frac{y_{1}}{1+\sigma^{2}/P}$ ;
/* Tx sends the error in estimate $\hat{\Theta}_{1}-\Theta$ over the next $D-1$ rounds */
while $2\leq i\leq D$ do
       Tx: Compute the error in estimate of previous round ${\epsilon}_{i-1}=\hat{\Theta}_{i-1}-\Theta$ ;
       Power normalization: $x_{i}=\frac{\sqrt{P}}{\sigma_{i-1}}{\epsilon}_{i-1}$, where $\sigma^{2}_{i-1}=\mathbb{E}[\epsilon^{2}_{i-1}]$;
       Forward channel: $y_{i}=x_{i}+n_{i}$;
       Rx: LMMSE estimate of transmit symbol $\hat{\epsilon}_{i-1}=\frac{\sqrt{P}\sigma_{i-1}}{P+\sigma^{2}}y_{i}$ ;
       Update the estimate of $\Theta$ as: $\hat{\Theta}_{i}=\hat{\Theta}_{i-1}-\hat{\epsilon}_{i-1}$
end while
Decoding: Map $\hat{\Theta}_{D}$ to the closest symbol in the $2^{K}$ PAM constellation.
@@
III-B Gallager-Nakibŏglu coding scheme
In [4], Gallager-Nakibŏglu proposed a two-phase scheme that builds on the Elias scheme, which is also closely related to the SK scheme. While the SK scheme considers the problem of transmitting a discrete symbol, Elias [27] studied the problem of transmitting a Gaussian random variable $U\in\mathcal{N}(0,\sigma^{2})$. The strategy for forward and feedback transmissions is similar to SK, where a scaled version of the error in the LMMSE estimate is transmitted in every round $i>1$. The main difference between the SK and Elias schemes lies in that the SK scheme aims to refine the message itself, while the Elias scheme aims to refine the noise added in the very first transmission.
GN scheme operates in two phases and relies on the assumption that after a sufficiently large number of rounds of the Elias scheme, the effective SNR for the forward channel shall be adequately high for the noise variance to be considered small compared to the distance between the symbols in the PAM constellation. In such a high-SNR regime, a strategy superior to the Elias scheme can be implemented by taking advantage of the discrete nature of the signal. Instead of transmitting the original error vector with respect to $2^{K}$-ary PAM, the integer difference in the PAM index of the estimate and the true PAM symbol is transmitted. We refer to this as discrete-symbol scheme. This method results in an error exponent that decreases with an exponential order, which increases linearly with the number of rounds. For this work, we assume that the high-SNR region is realized in the final round of communication, which is valid according to the 2-phase strategy described in [4].
We now describe the GN scheme in detail. The first round of GN is simply uncoded PAM, except for the power allocation. In [4], it is shown the optimal power distribution across $D$ rounds is attained by choosing $P_{1}$ and $P_{2}$ such that $P_{1}+(D-1)P_{2}=DP$ and $P_{1}=P_{2}+1$ where $P_{1}$ is the power constraint in round 1 and $P_{2}$ is the power constraint in remaining rounds. Hence, the transmission in round 1 is given by
@@ III-B Gallager-Nakibŏglu coding scheme
For the remaining rounds, the goal of the transmitter is to send the noise $n_{1}$ to the receiver as in the Elias scheme, where the LMMSE estimate at the receiver is improved iteratively, and a maximum likelihood (ML) detection is used at the end of $D-1$ rounds of communication to map the estimate to the original PAM constellation.
Finally, for the last round, GN uses a discrete-symbol scheme suitable for the high-SNR region by transmitting the error in the PAM index $U$. We follow the assumption from [4] that the high-SNR region guarantees that $U\in\{-1,0,1\}$ with high probability. In [4], an ML decoder is used for the ease of analysis across multiple rounds of discrete symbol scheme, which is sub-optimal. However, since we are assuming only one round of the discrete-symbol scheme in this work, we assume a maximum-a-posteriori (MAP) decoder, which is optimal for the performance. Hence, the final round of GN can be viewed as MAP detection on a constellation $\{-1,0,1\}$ with probability distribution $\{\nicefrac{{p_{GN1}}}{{2}},1-p_{GN1},\nicefrac{{p_{GN1}}}{{2}}\}$ where $p_{GN1}$ is the probability of error after $D-1$ rounds. The complete algorithm is described in Alg. 4 in Appendix Section A.
Error analysis.  The probability of error for a rate $\nicefrac{{K}}{{D}}$ Gallager-Nakibŏglu scheme can be computed in two phases. The first phase can be analyzed as 1 round of uncoded PAM followed by $D-2$ rounds of Elias scheme, for which it is shown in [4] that the probability of error is given by
@@
where $S$ is the SNR of the forward AWGN channel on a linear scale.
In the final round, the discrete integer difference between the PAM index of the decoded message $\hat{m}$ and the index of the true message $m$ is transmitted, where $(\hat{m}-m)\in\{-1,0,1\}$ with probability distribution $\{\nicefrac{{p_{GN1}}}{{2}},1-p_{GN1},\nicefrac{{p_{GN1}}}{{2}}\}$ based on the assumption from [4]. The probability of error in decoding using a MAP decoder can thus be computed as
@@
where $S$ is the SNR of the forward AWGN channel in linear scale and $\gamma$ is the detection threshold given by
$\gamma=\frac{1}{2\sqrt{p_{GN1}}}+\frac{\sqrt{p_{GN1}}}{S}\log\left(\frac{2(1-p_{GN1})}{p_{GN1}}\right),$ (12)
where $S$ is the SNR of the forward AWGN channel in linear scale.

### v3

@@ Abstract
The design of reliable and efficient codes for channels with feedback remains a longstanding challenge in communication theory. While significant improvements have been achieved by leveraging deep learning techniques, neural codes often suffer from high computational costs, a lack of interpretability, and limited practicality in resource-constrained settings. We focus on designing low-complexity coding schemes that are interpretable and more suitable for communication systems. We advance both analytical and neural codes. First, we demonstrate that PowerBlast, an analytical coding scheme inspired by Schalkwijk-Kailath (SK) and Gallager-Nakiboğlu (GN) schemes, achieves notable reliability improvements over both SK and GN schemes, outperforming neural codes in high signal-to-noise ratio (SNR) regions. Next, to enhance reliability in low-SNR regions, we propose LightCode, a lightweight neural code that achieves state-of-the-art reliability while using a fraction of memory and compute compared to existing deep-learning-based codes. Finally, we systematically analyze the learned codes, establishing connections between LightCode and PowerBlast, identifying components crucial for performance, and providing interpretation aided by linear regression analysis.
@@ I Introduction
Shannon introduced the concept of feedback channel in [1]. In a channel with feedback, the transmitter cooperates with the receiver to improve the probability of successful transmission by utilizing the feedback from the receiver. In [1], Shannon assumes a perfect noiseless feedback channel with unit delay and demonstrates that the availability of feedback at the transmitter does not change the capacity of the resultant forward channel for memoryless channels. Interestingly, while the capacity remains the same, significant improvements in error exponents can be achieved with the help of feedback.
One of the seminal schemes in the noiseless feedback setting has been provided by Shalkwijk and Kailath in [2, 3] (SK) using a simple linear encoding scheme resulting in a doubly exponential error exponent for finite block lengths. In [4], a two-phase variant of the SK scheme, which we refer to as Gallager-Nakiboğlu (GN), was discussed that can lead to further exponential decay of the error, provided a sufficiently good SNR is available for the forward channel. The SK scheme has been enhanced in various ways. The Modulo-SK scheme [5] and schemes by Chance-Love [6] and Mishra et al [7], extends the SK scheme for noisy feedback, while compressed error correction (CEC) [8] and accumulative iterative code (AIC) [9] focus on reducing channel use through continuous error vector compression, using noiseless feedback.
@@
Recent advances in deep learning revived the interest in coding for channels with feedback by leveraging the expressive power [12] of deep neural networks. Several works proposed deep learning approaches to improve the performance of channel codes, ranging from augmenting analytical decoders with learnable parameters [13, 14, 15, 16] to creating novel neural network architectures based neural encoders and decoders [17, 18] and improving the code design using sequential models [19, 20]. For channels with feedback, deep learning techniques were used to design new encoding and decoding schemes that take advantage of the high-capacity feedback channel. Deepcode [21] modeled both encoder and decoder functions as recurrent neural networks (RNNs) to process a bit stream in a sequential manner to directly minimize the end-to-end transmission error over an additive white Gaussian noise (AWGN) channel. Several works further explored this idea of using learning-based approaches for modeling the encoding and decoding operations, which can be broadly classified into the RNN family of codes [22, 23, 24] and the transformer family of codes [25, 26], discussed in detail in Section V. The current state-of-the-art is generalized block attention feedback (GBAF) [26], which uses self-attention and transformer architecture to perform block coding.
While these state-of-the-art deep-learning-based feedback codes provide tremendous improvements in BLER performance, they also come with significant memory and computational costs that may not be supported by the next generation of communication transceivers that operate with limited onboard resources. Therefore, it is desirable to develop lightweight codes that use simple schemes while providing desirable performance. To accomplish a reduced complexity coding scheme, we explore two different directions in this work. First, we review and understand the existing analytical feedback coding schemes to identify the limitations and propose a new feedback coding scheme that provides non-trivial performance improvements over the existing schemes for channels with passive, noiseless feedback. Next, we propose a lightweight deep-learning-based feedback coding scheme that can significantly reduce the complexity compared to neural block-feedback coding schemes by limiting ourselves to a symbol-by-symbol scheme instead of block coding schemes.
@@
  • We propose PowerBlast, an analytical coding scheme that noticeably improves the performance over Schalkwijk-Kailath and two-phase Gallager-Nakiboğlu schemes (Section IV) and exhibits reliability comparable to state-of-the-art deep-learning-based coding schemes in regions of high-SNR (Section VII).
@@
  • We analyze the representations learned by the LightCode encoder using linear regression and draw comparisons with PowerBlast. We also demonstrate that the relation between encoded symbols and the feedback is highly non-linear in the final rounds, underscoring the need for a deep-learning-based coding scheme (Section VIII).
@@ II System Model
We consider a feedback coding scheme with an AWGN forward channel and an AWGN feedback channel with noisy passive feedback. The goal is to transmit a message vector $\mathbf{u}\in\{0,1\}^{K}$ of length $K$ to the receiver using a total of $D$ independent channel uses i.e., the noise across the rounds is independent and identically distributed (i.i.d). This results in an overall coding rate of $R=\nicefrac{{K}}{{D}}$. In this work, we consider symbol-by-symbol coding, where the block of $K$ bits will be mapped to one symbol. Hence, the terms block and symbol can be used interchangeably, and the number of rounds of communication for a rate $\nicefrac{{K}}{{D}}$ code is $D$.
@@
where $\phi$ is the encoding function and $n_{1}\sim\mathcal{N}(0,\sigma^{2}_{ff})\in\mathbb{R}$ denotes the feedforward noise. The receiver then sends the noisy received symbol as feedback using the feedback channel as
@@
where $\psi$ is the decoding function and $\hat{\mathbf{u}}\in\{0,1\}^{K}$. The objective is to design an encoder-decoder pair $\{\phi,\psi\}$ that minimizes the probability of error $\mathbb{P}\{\mathbf{u}\neq\hat{\mathbf{u}}\}$ for a given number of rounds $D$, under a sum power constraint of $\sum_{i=1}^{D}\mathbb{E}[|x_{i}|^{2}]\ \leq\ D$.
@@ III Analytical Coding Schemes for Channels with Noiseless Feedback
In this section, we review analytical coding schemes for channels with noiseless feedback i.e., $\sigma_{fb}^{2}=0$. We begin by reviewing the celebrated Schalkwijk-Kailath coding scheme [2], one of the seminal works in coding for channels with noiseless feedback. We then review a less widely-known scheme by Gallager-Nakiboğlu [4]. This is similar to the SK scheme for the first $D-1$ rounds. Still, it deviates significantly in the last round, tailoring for the transmission of discrete messages, often significantly improving the performance in one round of communication. The Gallager-Nakiboğlu (GN) scheme is typically not considered as a baseline as it exhibits a worse error rate compared to SK in certain regions of SNR. However, in the next section, we introduce a novel analytical coding scheme, building on the SK and GN schemes. This new scheme achieves a significantly lower error rate than the SK or GN schemes and is often comparable to highly complex neural coding schemes in high-SNR regions, defying the conventional belief that neural coding schemes are much more reliable than analytical ones.
@@ III-A Schalkwijk-Kailath coding scheme
The SK scheme considers the problem of transmitting a fixed number of bits on an AWGN forward channel and a noiseless feedback channel. The transmission begins by mapping $K$ bits of information symbols to a single $2^{K}$-ary pulse-amplitude modulation (PAM) symbol $\Theta$ from the constellation
$\Theta\in\{\pm 1\eta,\pm 3\eta,\dots,\pm(2^{K}-1)\eta\},$ (6)
where $\eta=\sqrt{\nicefrac{{3}}{{(2^{2K}-1}})}$ is the scaling factor to ensure unit power normalization of the PAM constellation. Even though a block of information is transmitted since all the $K$ bits of information are mapped to one PAM symbol, the SK scheme at its core can be considered a symbol-by-symbol feedback scheme. We now describe the SK coding scheme in detail.
@@
where $n_{1}\in\mathcal{N}(0,\sigma_{ff}^{2})$ is the noise in the forward AWGN channel. The received symbol $y_{1}$ is then sent back to the transmitter noiselessly. For the second round, the transmitter first computes the receiver’s estimate of the transmitted symbol based on $y_{1}$ as $\hat{\Theta}_{1}=\frac{\sqrt{P}y_{1}}{P+\sigma^{2}_{ff}}$ using linear minimum mean-square error (LMMSE). In the second round, after receiving $y_{1}$ as feedback, the transmitter sends the scaled version of the error in the LMMSE estimate from the previous round, i.e., $\epsilon_{1}=\hat{\Theta}_{1}-\Theta$. This process continues for the remaining rounds. In other words, starting from round $i=2$, the goal of the transmitter is to transmit the error in estimate from the previous round, $\epsilon_{i-1}=\hat{\Theta}_{i-1}-\Theta$, after scaling appropriately to satisfy the power constraint. The complete algorithm is described in Algorithm 1.
@@
where $S$ denotes the SNR of the forward AWGN channel on a linear scale. Note that [5] assumes a minimum-variance unbiased estimator (MVUE) at the end of round $1$, $\hat{\Theta}_{1}=\frac{y_{1}}{\sqrt{P}}$, for ease of analysis but it is sub-optimal in terms of error in the estimate after round 1.
1
Input: Message symbol $\Theta$, number of rounds $D$, forward noise variance $\sigma_{ff}^{2}$
2
3 Round 1: Tx: Power normalization: $x_{1}=\sqrt{P}\Theta$;
4 Forward channel: $y_{1}=x_{1}+n_{1}$;
5 Rx: LMMSE estimate of transmit symbol $\hat{\Theta}_{1}=\frac{\sqrt{P}y_{1}}{P+\sigma^{2}_{ff}};$
6 /* Tx communicates the error in estimate $\hat{\Theta}_{1}-\Theta$ over the next $D-1$ rounds */
7
8while $2\leq i\leq D$ do
9       Tx: Compute the error in estimate of previous round ${\epsilon}_{i-1}=\hat{\Theta}_{i-1}-\Theta$ ;
10       Power normalization: $x_{i}=\frac{\sqrt{P}}{\sigma_{i-1}}{\epsilon}_{i-1}$, $\sigma^{2}_{i-1}=\mathbb{E}[\epsilon^{2}_{i-1}]$;
11       Forward channel: $y_{i}=x_{i}+n_{i}$;
12       Rx: LMMSE estimate of transmit symbol $\hat{\epsilon}_{i-1}=\frac{\sqrt{P}\sigma_{i-1}}{P+\sigma_{ff}^{2}}y_{i}$ ;
13       Update the estimate of $\Theta$ as: $\hat{\Theta}_{i}=\hat{\Theta}_{i-1}-\hat{\epsilon}_{i-1}$
14 Decoding: Map $\hat{\Theta}_{D}$ to the closest symbol in the $2^{K}$ PAM constellation.
@@ 13       Update the estimate of $\Theta$ as: $\hat{\Theta}_{i}=\hat{\Theta}_{i-1}-\hat{\epsilon}_{i-1}$
III-B Gallager-Nakiboğlu coding scheme
In [4], Gallager-Nakiboğlu proposed a two-phase scheme that builds on the Elias scheme [27], also closely related to the SK scheme. While the SK scheme considers the problem of transmitting a discrete symbol, Elias studied the problem of transmitting a Gaussian random variable $U\sim\mathcal{N}(0,\sigma^{2})$. The strategy for forward and feedback transmissions is similar to SK, where a scaled version of the error in the LMMSE estimate is transmitted in every round $I>1$. The main difference between the SK and Elias schemes lies in that the SK scheme aims to refine the message itself, while the Elias scheme aims to refine the estimate of noise added in the very first transmission.
GN scheme operates in two phases and relies on the assumption that after a sufficiently large number of rounds of the Elias scheme, the effective SNR for the forward channel shall be adequate for the noise variance to be considered small compared to the distance between the symbols in the PAM constellation. In such a high-SNR regime, a strategy superior to the Elias scheme can be implemented by taking advantage of the discrete nature of the signal. Instead of transmitting the original error vector with respect to $2^{K}$-ary PAM, the integer difference between the PAM index of the estimate and the true PAM symbol is transmitted. We refer to this as discrete-symbol scheme. This method results in an error exponent that decreases with an exponential order, which increases linearly with the number of rounds. For this work, we assume that the high-SNR region is realized in the final round of communication, which is valid according to the 2-phase strategy described in [4].
We now describe the GN scheme in detail. The first round of GN is simply uncoded PAM, except for the power allocation. In [4], it is shown the optimal power distribution across $D$ rounds is attained by choosing $P_{1}$ and $P_{2}$ such that $P_{1}+(D-1)P_{2}=DP$ and $P_{1}=P_{2}+1$, $P_{1}$ is the power constraint in round 1 and $P_{2}$ is the power constraint in remaining rounds. Hence, the transmission in round 1 is given by
@@ III-B Gallager-Nakiboğlu coding scheme
For the remaining rounds, the goal of the transmitter is to communicate the noise $n_{1}$ to the receiver as in the Elias scheme, where the LMMSE estimate at the receiver is improved iteratively, and a maximum likelihood (ML) detection is used at the end of $D-1$ rounds of communication to map the estimate to the original PAM constellation.
Finally, for the last round, GN uses a discrete-symbol scheme suitable for the high-SNR region by transmitting the error in the PAM index $U$. We follow the assumption from [4] that the high-SNR region guarantees that $U\in\{-1,0,1\}$ with high probability. In [4], an ML decoder is used for the ease of analysis across multiple rounds of discrete symbol schemes, which is sub-optimal. However, since we assume only one round of the discrete-symbol scheme in this work, we assume a maximum-a-posteriori (MAP) decoder, which is optimal for the performance. Hence, the final round of GN can be viewed as MAP detection on a constellation $\{-1,0,1\}$ with probability distribution $\{\nicefrac{{p_{GN1}}}{{2}},1-p_{GN1},\nicefrac{{p_{GN1}}}{{2}}\}$, $p_{GN1}$ is the probability of error after $D-1$ rounds. The complete algorithm is described in Algorithm 4 in Appendix, Section C.
Error analysis.  The probability of error for a rate $\nicefrac{{K}}{{D}}$ Gallager-Nakiboğlu scheme can be computed in two phases. The first phase can be analyzed as one round of uncoded PAM followed by $D-2$ rounds of Elias scheme, for which it is shown in [4] that the probability of error is given by
@@
where $S$ denotes the SNR of the forward AWGN channel on a linear scale.
In the second phase, which corresponds to the final round, the discrete integer difference between the PAM index of the decoded message $\hat{M}$ and the index of the true message $M$ is transmitted. Here, the message index $M\in\{0,1,2,\dots,2^{K}-1\}$ is deterministic based on the transmitted symbol $\Theta\in\{\pm 1\eta,\pm 3\eta,\dots,\pm(2^{K}-1)\eta\}$ and can be computed using a predetermined mapping, where $\eta$ is the scaling factor to ensure unit power normalization. Similarly, $\hat{M}$ corresponds to the message index whose corresponding symbol is closest to the estimated transmitted symbol $\hat{\Theta}$.
As the error in the $(\hat{M}-M)$ lies in $\{-1,0,1\}$ with probability distribution $\{\nicefrac{{p_{GN1}}}{{2}},1-p_{GN1},\nicefrac{{p_{GN1}}}{{2}}\}$ based on the assumption from [4], the probability of error in decoding using a MAP decoder can be computed as
@@
where $S$ is the SNR of the forward AWGN channel on a linear scale, and $\gamma$ is the detection threshold given by
$\gamma=\frac{1}{2\sqrt{p_{GN1}}}+\frac{\sqrt{p_{GN1}}}{S}\log\left(\frac{2(1-p_{GN1})}{p_{GN1}}\right).$ (12)